OpenStudy (anonymous):

A sign designer created a sign in the shape shown below. The semicircles are congruent to one another with a diameter of 4 inches. What is the perimeter of the sign? Answer about 38 inches about 78 inches about 76 inches about 80 inches

OpenStudy (anonymous):

This is the sign:

OpenStudy (anonymous):

The perimeter is the sum of the 4 sides. You know that two sides are 20 inches long, but the other two sides are a bit tricky since they are made up of 3 semicircles each. Since the circumference (perimeter) of a circle is \[\Pi(d)\] Where d is the diameter. Then half a circle's (semicircle) circumference (perimeter) must be \[1/2\Pi(d)\] Letting \[\Pi=3.14\] The perimeter of one of the semicircles must be \[1/2*3.14*4=6.28\] So the perimeter of three semicircles must be 3*6.28=18.84 inches So now it's just the sum of all four sides. 18.84 inches + 20 inches + 18.84 inches + 20 inches = 77.68 inches